XAMPLE 28-2 Force between two current carrying wires. The two wires of a 2.0-m-long appliance cord are 3.0 mm apart and carry a current of O A de. Calculate the force between these wires
Answer #1

solution

(a)

Givne that,

L = 2 m

d = 3 mm

l1 = l2 = 8 A

Force b/w thses wires,

F = u0*l1*l2*L / 2*pi*d

F = 4*pi*10-7*8*8*2 / 2*pi*0.003

F = 8.53*10-3 N

i hope it will be helpful for you
Kindly Upvote if Helpful :)